OpenStudy (anonymous):

GIVING MEDAL! PLEASE HELP! For f(x)=4x+1 and g(x)=x^2-5, find (f+g)(x) A. x^2+4x-4 B. 4x^2-19 C. x^2+4x+6 D. 4x^3-4

OpenStudy (anonymous):

Answer is "A"

OpenStudy (praxer):

add both the functions as the theory of simple algebra follows and then we shall find that, f+g(x)= f(x)+g(x) =4x+1+x^2-5 =x^2+4x+(1-5) =x^2+4x-4 did you get it @yara26